Please help me find the inverse of the function: +y=+6+log+base+5+x+
Answer by lwsshak3(11628)

Please help me find the inverse of the function: y= 6 log base 5 x
The inverse of a log function is the exponential function:
(base raised to the log of a number=number)
..
For given equation:
y= 6 log base 5 x
y= log base 5 (x^6)
..
base=5
log of number=y
number=x^6
..
Inverse: 5^y=x^6